agnose U lamphenicol b. Enrofloxacin c. Optimmune d. Acarexx 21. How many milligrams of tropicamide are in a 0.5% solution? a could you please check the answers for 21. 22. 23 and 24. please. I would like to see if I got these right
Answer #1

Ans 21: 5 mg

The solution's mass by volume% cocentration, m/v %, tells the number of grams of solute present for every 100 ml of the solution .

Therefore 0.5 % solution = 0.5 / 100 = 0.005 g solute

convert it to mg = 0.005 * 1000 = 5 mg ( as 1gm = 1000 mg)

Ans 22 : 16.8 ml

Veterinarian prescribed 0.15 % solution.

0.15 % solution means= 0.15 g solute in 100 ml solution = 0.15 / 100 = 0.0015 = 1.5 g ( multiplied by 1000 to convert in mg )

This solution needs to be adminnistered every 4 hours in a day.

1 day has total 24 hours .Therefore patient is getting total 6 time dose in a day .( 24 / 4 = 6 )

Therefore for 14 days he will get total 84 doses.

Single dose contains volume of 0.2 ml .

Therefore for 84 doses it will be 16.8 ml. ( 84 * 0.2 = 16.8 )

Ans 23 : 7.5 drops

If 2drops are required to anesthetize cornea for 8 minutes

Then X drops are required for half hour ( i.e 30 min)

cross multiplication;

2 * 30 = 8 * X

X=2 * 30 / 8 = 60 / 8 = 7.5 drops

Ans 24 : 84 strips

If 21 dogs are stained by fluoresecin strips in a week .

Then 21 strips are required .( 1 for each )

There are 4 weeks in a month.

If 21 strips for 1 week .

Then for 4 weeks =21 *4 = 84 strips